Laplace transform of the unit step functionY + 3y + 2y = uz(t).
Answer #1

Given equation is

y + 3y + 2y = uz(t)

Taking Laplace transform of both sides, we get

L{y} +3L{y} + 2L{y} = L*uz(t)}

25 e s?Y(s) – sy(0) - y0) + 3(sY (s) – 4 (0)) + 2Y(s) = S25 e $Y(s) – sy(0) – }(0) + 3sY(s) – 3y (0) + 2Y(s) = S

- 28 using given conditions (8²+38+2) Y18) - As-B-3A (8+2)(8+1) 718) A8+ (3A+B) + é 8 -28 8 -28 +(3A+B) at 8t1 b &+2 + (3A+B)